A piece of beef tendon 25 cm long and of cross section 7.8mm*13 mm is stretched to a length of 25.19 cm when it supported a vertical load of .96kg. Calculate young's modulus for the tendon
Answer #1

L = 25 cm = 0.25 m
A = 7.8*13*10^-6=1.014*10^-4 m²
F = mg=0.96*9.8= 9.408 N
e = 0.19 cm = 0.0019 m

Now young's modulus E = FL/(eA)
. .= 9.408 x 0.25 / (0.0019 x 1.014 x10^-4)
. .= 12.2 x10^6 Pa =12.2 MPa